Watch the moon!

We know that the moon shines by the reflected light from the Sun. Now the Sun always produces the same amount of sunshine. But the Sun is not always the same distance away from the Earth/Moon. Right now, the Sun is closest to us (see the article on Solstices for details). Because of this, more sunlight reaches us and so the moon looks brighter than usual. In fact, the brightest moon would have been on the full-moon day closest to Dec 21, 2008 (southern solstice). The moon will still be bright for the next few weeks. By the time it is June, the Sun is farthest from us, and the moon is the dullest. After that, it will again get brighter till the December solstice is again reached.

So be sure and keep on looking at the moon every full-moon day, all year. Watch it especially in the evening, at moon-rise time. If you are lucky enough to live near the (eastern) sea, you will see some spectacular moons.